Deshaun invested his savings in two investment funds. The amount he invested in Fund A was 3 times as much as the amount he invested in Fund B. Fund A returned a 5% profit and Fund B returned a 8% profit. How much did he invest in Fund B, if the total profit from the two funds together was $1610?
Answer by richwmiller(17219)

Total yearly interest for the two accounts is: $1610
x=3y+0
0.05*x+0.08*y=1610
Substitute for x
0.05*(3y+0)+0.08*y=1610
Multiply out
0.15y+0+0.08*y=1610
Combine like terms.
0.15y+0+0.08*y=1610
0.15y+0.08*y=1610-0
0.23*y=1610
Isolate y
0.23*y=1610
y=1610/0.23
y=$7000 at 8% Fund B
y rounded off
y=$ 7000.00 at 8% Fund B
x=3y+0
x=3*7000+0
x=21000+0
x=$21000 at 5%
x rounded off
x=$ 21000.00 at 5% Fund A
x=$21000.00 at 5% Fund A
y=$ 7000.00 at 8% Fund B
Check
0.05*21000+0.08*7000=1610
1050+560=1610
1610=1610
If this statement is TRUE and neither x nor y is negative then all is well
codeinttm